how would you solve an equation set up with X as a dinominator? such as two over X = 3?
Found 2 solutions by Fombitz, nyc_function:
Answer by Fombitz(32388)

Mutliply both sides of the equation by X.
Then divide by 3.

Answer by nyc_function(2741)

2/x = 3
Multiply both sides of the equation by x to remove the fraction.

(2/x) * x = 3 * x
2 = 3x
Now divide both sides by 3 to find x.

2/3 = x

Understand?